# Schrödinger Bridge Matching for Tree-Structured Costs and Entropic Wasserstein Barycentres

### v1

@@ C.1 Implementation Details and Design Choices
We implement the TreeDSBM procedure using the JAX framework [Bradbury et al., 2018]. Below, we outline some of the design considerations for implementing the method.
@@
In our implementation, we use separate neural networks to parameterise the vector fields for each direction along each edge, totalling $|\mathcal{E}|$ networks in total. Note that one could alternatively use a shared network along each edge with an additional binary input indicating the direction; this parameterisation was used in De Bortoli et al. [2024].
@@
As we train both directions along each edge, simulation of the SDEs along the tree structure can be initialised from any of the observed nodes in $\mathcal{S}$. Following Shi et al. [2023], one can rotate the starting vertex between IMF iterations. This helps to mitigate any drift that accumulates in the marginals, as the coupling samples will only use true samples from the current starting marginal; for an analysis of this, see De Bortoli et al. [2024]. In our experiments, we often did not observe noticeable drift in the marginals, and instead generated samples in the coupling by initialising equally across the observed vertices in $\mathcal{S}$. In our sampling, we used an Euler-Maruyama discretisation scheme with 50 uniformly-spaced steps.
@@
Other than image experiments, we use a basic MLP-based vector field model. It consists of an MLP spatial embedding with hidden layers [128, 256] to embed into dimension 32, a time embedding consisting of a sine positional-encoding and an MLP with hidden layers [128, 256] also embedding into dimension 32, before concatenating the embeddings and passing through an MLP with hidden layers [512, 256, 128]. This is the same architecture used in De Bortoli et al. [2021], Noble et al. [2023]. In image experiments, we use the UNet architecture [Ronneberger et al., 2015] with the improvements from Dhariwal and Nichol [2021], using the JAX implementation from Song et al. [2023]. For all experiments, we use the Adam optimiser [Kingma and Ba, 2015] with default parameters 0.9 and 0.999.
@@
We have presented the methodology according to using Brownian reference measures along each edge. However, the methodology extends to other reference process, as long as the respective bridging processes conditioned on the values at $\mathcal{S}$ are tractable (for example, this is true for Ornstein-Uhlenbeck processes). The connection to quadratic-cost optimal transport is however less simple beyond the Brownian case. See Shi et al. [2023] for a more detailed treatment of this general case.
@@ C.2 Methodology Extensions for Improving Convergence Speed
The TreeIMF procedure can be initialised with any coupling $\Pi_{\mathcal{S}}^{0}$ over $\mathcal{S}$ with correct marginals. For the standard SB problem, Tong et al. [2024b] note that the SB solution is a mixture of bridges mixed by a static $\varepsilon$-OT solution, and thus propose a single iteration of bridge matching on samples generated by a static $\varepsilon$-OT solver applied on minibatches (see also Pooladian et al. [2023], Tong et al. [2024a], Fatras et al. [2021]).
Similarly, one can initialise TreeIMF using samples obtained from static mmOT solvers applied to minibatches. Such a procedure can speed up convergence to the TreeSB solution by initialising closer to the true solution. We remark, however, that such minibatching approaches can incur large errors (particularly in higher dimensions or for small minibatches; for example the Wasserstein-1 error grows as $O(B^{-1/(2d)})$ [Sommerfeld et al., 2019]), and so the advantages of such methods are lessened as dimensionality increases.
@@
Iterative Markovian Fitting presents a mathematically elegant approach for solving the SB problem, with significant practical improvements over the IPF procedure. However, the iterative nature of the algorithm remains a downside - each iteration of the two-step procedure involves first simulating the current Markovian process, and then retraining a neural network with the bridge-matching loss. In practice, one might wonder if it is possible to perform the simulations and bridge-matching procedure simultaneously to avoid the expensive iterative nature of the algorithm. Recent work De Bortoli et al. [2024] answered this in the affirmative and propose the $\alpha$-IMF procedure, which instead corresponds to a discretisation of a continuous flow of processes that converge to the SB. One could naturally design an analogous methodology for the tree-based setting that we consider; we leave such extensions for future work.
@@ Appendix D Experimental Details
[…]
@@
We follow the experimental setup of Noble et al. [2023]. The marginals consist of moon, circle, and spiral datasets from scikit-learn [Pedregosa et al., 2011], centred and scaled by a factor of 7.0. We aim to learn the ($\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3}$)-barycentre of the dataset.
@@
For the ground-truth, we use the in-sample free-support barycentre algorithm of Cuturi and Doucet [2014] implemented in Flamary et al. [2021], with 1500 datapoints. Note that the aim of this experiment is not to outperform in-sample methods; it is known that such approaches perform well in low dimensions, but do not scale well as dimension increases. The in-sample method is used here to provide a close approximation to the ground-truth, allowing us to judge the success of the continuous Wasserstein-2 barycentre approaches that we compare.
For TreeDSB, we use the checkpoints provided by Noble et al. [2023] which were trained for 50 IPF iterations.
We also report results for the WIN algorithm from Korotin et al. [2022]. This is an iterative algorithm inspired by Álvarez Esteban et al. [2016]; it pushes forward a source latent distribution $\rho$ through a function $G$ to give a generative model $\nu=G\#\rho$ for the barycentre, and also learns maps $T_{i}$ and $T_{i}^{-1}$ transporting from the generated barycentre to the marginals and vice versa. In our experiments, the barycentre generator $\nu=G\#\rho$ was unable to fit the true barycentre accurately, nor were the maps $T_{i}^{-1}\#\mu_{i}$. We hypothesise that the neural maps struggle to model the discontinuous transports well. However, the combined map $(\sum_{i}\lambda_{i}T_{i})\#\nu$ was able to give a good approximation of the true barycentre, which are the results we report in Figure 2 and Table 2.
Additionally, we applied the W2CB [Korotin et al., 2021] and NOTWB [Kolesov et al., 2024a] algorithms to this example, but were unable find hyperparameters to make the algorithms to converge to the correct solution. Again, this may be due to the difficulty in modelling the discontinuous transports with neural networks. We also anticipate that the loss landscapes caused by this complex example may have caused these methods to get stuck in local minima. It appears that the iterative schemes of TreeDSBM, TreeDSB, and WIN aid in overcoming such issues.
@@
Our TreeDSBM implementation took approximately 2 minutes for each IMF iteration when training the edges sequentially, and took around 15 minutes to run the 6 IMF iterations. Note that this could be improved by a factor of around 3 by training the edges simultaneously, and one could also obtain good results using fewer IMF iterations.
@@
Additional Plots
In Figure 6, we plot TreeDSBM samples obtained for different values of entropy-regularisation $\varepsilon$. As expected, increasing $\varepsilon$ leads to a slight blurring bias in the solution.
In Figure 6, we demonstrate the progression of the TreeDSBM barycentre approximation as we run the IMF iterations. We run with $\varepsilon=0.1$, and plot samples generated from leaf 00. We see that even after only two IMF iterations, TreeDSBM gives a good approximation to the solution. Such behaviour reflects similar results reported in von Lindheim [2023], which observes that iterative fixed-point approaches [Álvarez Esteban et al., 2016] exhibit very fast convergence to the solution. TreeDSBM performs only a single bridge-matching iteration along each edge before updating the barycentre, rather than computing full OT maps (which would be expensive). As such, it strikes a good balance between the efficiency of iterative fixed-point-based approaches, without requiring full OT map computations before updating the barycentre approximation.
@@ D.2 MNIST 2,4,6 Barycentre
(a) Barycentre samples, $\varepsilon=0.01$.
(b) Samples as $\sum_{i}\lambda_{i}Y_{i}$, $\varepsilon=0.01$.
@@
We also follow the experimental setup of Noble et al. [2023] and compute the ($\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3}$)-barycentre of the 2,4 and 6 digits in the MNIST dataset [LeCun et al., 2010].
@@
We use the UNet architecture of Song et al. [2023], with 64 channels, channel multiples of (1,2,2), attention at layers (16,8), and 2 residual blocks at each layer. We train each bridge-matching procedure with 10,000 steps with batch size 64, at a learning rate of 1e-4 and with exponential moving average weight of 0.999. For subsequent IMF iterations, we simulate 8,192 coupling samples. We run for 4 IMF iterations, beyond which we did not see much change between iterations. The TreeDSB samples plotted in Figure 3 are those displayed in Noble et al. [2023], as we did not have the computational resources to run TreeDSB to convergence in this setting.
@@
In Figure 3, the TreeDSBM barycentre has some noise present in the samples. We remark here that this noise should be present in an accurately computed solution, as we are solving for the $\varepsilon$-TreeSB solution which adds entropic regularisation. If one wishes for less noise in the samples, one can run with a smaller entropy-regularisation value $\varepsilon$. Alternatively, one could generate samples by sampling $Y_{i}$ according to the learned coupling, and weighting them as $\sum_{i}\lambda_{i}Y_{i}$ (though note that would not be clear what kind of barycentre such samples would be from). We plot results for a smaller $\varepsilon=0.001$, and for the weighted coupling samples $\sum_{i}\lambda_{i}Y_{i}$ for $\varepsilon=0.01$ in Figure 7, to demonstrate that TreeDSBM can generate samples with minimal noise if so desired.
@@ D.3 Subset Posterior Aggregation
Hyperparameters:
@@
For the alternative methods, we use the implementations in their publicly available code (to which we provide the links in Appendix F), and we are grateful for their easy-to-use examples provided. For W2CB, we use a learning rate of $1\mathrm{e}{-4}$ in the negative binomial setting. For WIN, in the negative binomial case we rescale the source $z$-sampler by a factor of 10.0 to match the scale of the data better; without this, it did not appear to converge. We run W2CB and WIN for 10000 training iterations, and NOTWB for 2500 iterations. Other than those mentioned, we use the default parameters provided in the respective codebases.
@@
The W2CB algorithm appears to give good results after approximately 1000 training steps in both cases, which took around 10 minutes in our experiments. WIN converges after around 2500 iterations, which took around 45 minutes. NOTWB converged quickly after only around 400 iterations, which took around 5 minutes.
In both experiments, each TreeIMF iteration for our TreeDSBM implementation took approximately 1 minute when training the edges sequentially. TreeDSBM converged well using 4 IMF iterations, so training took around 5 minutes (including time for simulating training samples for the next iteration). This is competitive with the fastest alternative method NOTWB. Moreover, note that if the edges were trained simultaneously, this could be sped up further by a factor of around 5.
D.4 Higher-dim Gaussian Experiments

### v2

@@ C.1 Implementation details and design choices
    We implement the TreeDSBM procedure using the JAX framework [9]. Below, we outline some of the design considerations for implementing the method.
@@
[…]
@@
    As we train both directions along each edge, simulation of the SDEs along the tree structure can be initialised from any of the observed nodes in $\mathcal{S}$. Following [77], one can rotate the starting vertex between IMF iterations. This helps to mitigate any drift that accumulates in the marginals, as the coupling samples will only use true samples from the current starting marginal; for an analysis of this, see [23]. In our experiments, we often did not observe noticeable drift in the marginals, and instead generated samples in the coupling by initialising equally across the observed vertices in $\mathcal{S}$. In our sampling, we used an Euler-Maruyama discretisation scheme with 50 uniformly-spaced steps.
@@
    Other than image experiments, we use a basic MLP-based vector field model. It consists of an MLP spatial embedding with hidden layers [128, 256] to embed into dimension 32, a time embedding consisting of a sine positional-encoding and an MLP with hidden layers [128, 256] also embedding into dimension 32, before concatenating the embeddings and passing through an MLP with hidden layers [512, 256, 128]. This is the same architecture used in [24, 64]. In image experiments, we use the UNet architecture [72] with the improvements from [26], using the JAX implementation from [82]. For all experiments, we use the Adam optimiser [44] with default parameters 0.9 and 0.999.
    We note that for pointcloud experiments, capacity of the neural networks is not a limiting factor—any sufficiently-expressive network will work similarly well and fairly small MLP networks suffice, so using several networks for the different edges does not pose issues. One could also use a single network across the edges and additionally condition on the edge. This makes particular sense for problems with shared structure between edges, such as those for image data (and indeed these are settings for which the networks would be larger, and maintaining multiple networks could become a computational bottleneck).
    Figure 4: The two stages of the TreeIMF procedure, for a non-star-shaped tree structure. On this tree, the marginals at the leaf vertices $\mathcal{S}=\{1,3,5,6,7\}$ (blue) are fixed. The marginals at vertices $\mathcal{V}\backslash\mathcal{S}$ (red) are not fixed, and change during the procedure.
    Memory requirements
    The parallel nature of the TreeDSBM algorithm during training provides practitioners with a trade-off between memory consumption and wall-clock time. Namely, one can train the edges simultaneously if compute allows (either on a single GPU if enough memory, or parallelised across GPUs). If this cannot be done, one can train sequentially instead (in which case memory requirements for training would be comparable with standard bridge-matching). We report GPU consumption for different experiments in Table 5 (for sequential and joint training), for the hyperparameters used in the paper. These can of course be changed significantly by changing hyperparameters such as batch size. We note that in some of these experiments (e.g. the Gaussian experiments), peak GPU usage is due to the simulation and storage of the training samples for subsequent IMF steps, rather than during the network training. This can be reduced significantly by simulating in smaller batches, or by updating the cache during training rather than simulating all beforehand.
    We remark that TreeDSBM has improved memory requirements compared to TreeDSB. In TreeDSB, training a time-reversal along an edge requires saving entire trajectories simulated along the reverse direction, whereas TreeDSBM only requires storing endpoints (this is one of the most significant benefits of IMF over IPF).
    Table 5: Comparing memory usage (in MB) for sequential and joint training, for the hyperparameters used.
    Sequential Training (MB) Joint Training (MB)
    2d 447 687
    Data Aggregation (Poisson) 705 1219
    Gaussian ($d=64$) 1239 1239
    MNIST 2,4,6 4683 6407
@@
    We have presented the methodology according to using Brownian reference measures along each edge. However, the methodology extends to other reference process, as long as the respective bridging processes conditioned on the values at $\mathcal{S}$ are tractable (for example, this is true for Ornstein-Uhlenbeck processes). The connection to quadratic-cost optimal transport is however less simple beyond the Brownian case. See [77] for a more detailed treatment of this general case.
@@ C.2 Methodology extensions for improving convergence speed
    The TreeIMF procedure can be initialised with any coupling $\Pi_{\mathcal{S}}^{0}$ over $\mathcal{S}$ with correct marginals. For the standard SB problem, [88] note that the SB solution is a mixture of bridges mixed by a static $\varepsilon$-OT solution, and thus propose a single iteration of bridge matching on samples generated by a static $\varepsilon$-OT solver applied on minibatches (see also [70, 87, 30]).
    Similarly, one can initialise TreeIMF using samples obtained from static mmOT solvers applied to minibatches. Such a procedure can speed up convergence to the TreeSB solution by initialising closer to the true solution. We remark, however, that such minibatching approaches can incur large errors (particularly in higher dimensions or for small minibatches; for example the Wasserstein-1 error grows as $O(B^{-1/(2d)})$ [81]), and so the advantages of such methods are lessened as dimensionality increases.
@@
    Iterative Markovian Fitting presents a mathematically elegant approach for solving the SB problem, with significant practical improvements over the IPF procedure. However, the iterative nature of the algorithm remains a downside - each iteration of the two-step procedure involves first simulating the current Markovian process, and then retraining a neural network with the bridge-matching loss. In practice, one might wonder if it is possible to perform the simulations and bridge-matching procedure simultaneously to avoid the expensive iterative nature of the algorithm. The recent work of [23] answered this in the affirmative and propose the $\alpha$-IMF procedure, which instead corresponds to a discretisation of a continuous flow of processes that converge to the SB. We anticipate one could design an analogous methodology for the tree-based setting that we consider; we leave such extensions for future work.
@@ Appendix D Experimental details
    D.1 Synthetic 2dd barycentre
    Figure 5: TreeDSBM samples in the 2dd experiment, comparing different regularisation values $\varepsilon$.
@@ D.1 Synthetic 2dd barycentre
    We follow the experimental setup of [64]. The marginals consist of moon, circle, and spiral datasets from scikit-learn [66], centred and scaled by a factor of 7.0. We aim to learn the ($\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3}$)-barycentre of the dataset. This is a challenging problem—the lower-dimensional and discontinuous structures in the marginals mean that the barycentre has a complex and fragmented support, and the transport maps to the barycentre are highly discontinuous.
@@
    For the ground-truth, we use the in-sample free-support barycentre algorithm of [18] implemented in Python Optimal Transport [32], with 1500 datapoints. Note that the aim of this experiment is not to outperform in-sample methods; it is known that such approaches perform well in low dimensions, but do not scale well as dimension increases. The in-sample method is used here to provide a close approximation to the ground-truth, allowing us to judge the success of the continuous Wasserstein-2 barycentre approaches that we compare.
    For TreeDSB, we use the checkpoints provided by [64] which were trained for 50 IPF iterations.
    We also report results for the WIN algorithm from [47]. This is an iterative algorithm inspired by [4]; it pushes forward a source latent distribution $\rho$ through a function $G$ to give a generative model $\nu=G\#\rho$ for the barycentre, and also learns maps $T_{i}$ and $T_{i}^{-1}$ transporting from the generated barycentre to and from the marginals respectively. In our experiments, the barycentre generator $\nu=G\#\rho$ was unable to fit the true barycentre accurately, nor were the maps $T_{i}^{-1}\#\mu_{i}$. We hypothesise that the neural maps struggle to model the discontinuous transports well. However, the combined map $(\sum_{i}\lambda_{i}T_{i})\#\nu$ was able to give a good approximation of the true barycentre, which are the results we report in Figure 2 and Figure 2.
    Additionally, we applied the W2CB [49] and NOTWB [45] algorithms to this example, but were unable find hyperparameters to make the algorithms to converge to the correct solution. Again, this may be due to the difficulty in modelling the discontinuous transports with neural networks. We also anticipate that the loss landscapes caused by this complex example may have caused these methods to get stuck in local minima. It appears that the iterative schemes of TreeDSBM, TreeDSB, and WIN aid in overcoming such issues.
    Convergence speed
    In Figure 6, we demonstrate the progression of the TreeDSBM barycentre approximation as we run the IMF iterations. We run with $\varepsilon=0.1$, and plot samples generated from leaf 0. In Table 6, we also show how the Sinkhorn divergence to the ground truth evolves as the IMF iterations progress. After only two IMF iterations, TreeDSBM already gives a good approximation to the solution. Such behaviour reflects similar results reported in [55], which observes that iterative fixed-point approaches [4] exhibit very fast convergence to the solution. TreeDSBM performs only a single bridge-matching iteration along each edge before updating the barycentre, rather than computing full OT maps (which would be expensive). As such, it strikes a good balance between the efficiency of iterative fixed-point-based approaches, without requiring full OT map computations before updating the barycentre approximation.
@@
    Our TreeDSBM implementation took approximately 1 minute for each IMF iteration when training the edges jointly, and took around 7 minutes to run the 6 IMF iterations. Note that one could also obtain good results using fewer IMF iterations or fewer training steps.
@@
    Additional results
    In Figure 6, we plot TreeDSBM samples obtained for different values of entropy-regularisation $\varepsilon$. As expected, increasing $\varepsilon$ leads to a slight blurring bias in the solution. In Table 6 we also report the quality of fit to the marginals, calculated by simulating from the ‘moon’ marginal to the centre and then out to the other leaf nodes, and averaging the resulting Sinkhorn divergences to ground-truth samples from these marginals. We see that in this experiment there is negligible drift accumulation in the marginals.
@@ D.2 MNIST 2,4,6 barycentre
    (a) Barycentre samples, $\varepsilon=0.02$.
    (b) Samples as $\sum_{i}\lambda_{i}Y_{i}$, $\varepsilon=0.02$.
@@
    We also follow the experimental setup of [64] and compute the ($\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3}$)-barycentre of the 2,4 and 6 digits in the MNIST dataset [52].
@@
    We use the UNet architecture of [82], with 64 channels, channel multiples of (1,2,2), attention at layers (16,8), and 2 residual blocks at each layer. We train each bridge-matching procedure with 10,000 steps with batch size 64, at a learning rate of 1e-4 and with exponential moving average weight of 0.999. For subsequent IMF iterations, we simulate 8,192 coupling samples. We run for 4 IMF iterations, beyond which we did not see much change between iterations. The TreeDSB samples plotted in Figure 3 are those displayed in [64], as we did not have the computational resources to run TreeDSB to convergence in this setting.
@@
    In Figure 3, the TreeDSBM barycentre has some noise present in the samples. We remark here that this noise should be present in an accurately computed solution, as we are solving for the $\varepsilon$-TreeSB solution which adds entropic regularisation. If one wishes for less noise in the samples, one can run with a smaller entropy-regularisation value $\varepsilon$. Alternatively, one could generate samples by sampling $Y_{i}$ according to the learned coupling, and weighting them as $\sum_{i}\lambda_{i}Y_{i}$ (though note that would not be clear what kind of barycentre such samples would be from). We plot results for a smaller $\varepsilon=0.001$, and for the weighted coupling samples $\sum_{i}\lambda_{i}Y_{i}$ for $\varepsilon=0.02$ in Figure 7, to demonstrate that TreeDSBM can generate samples with minimal noise if so desired.
    Convergence speed
    We provide an analysis of the convergence speed of TreeDSBM in terms of the number of IMF iterations required. Unfortunately there is no ground-truth barycentre to compare to in this example; this makes quantitative evaluation of the obtained barycentre difficult. We therefore instead assess the fit to the marginals (from which we have true samples) as a proxy for the success of the algorithm, along with the transport cost — the transport cost provides an indication of the optimality of the maps, while if the marginals are not fitted accurately, then the resulting barycentre will be unreliable. While not an ideal measure of the ‘quality’ of the barycentre itself, this does provide a quantitative and, importantly, tractable proxy for the ‘success’ of the algorithm.
    To this end, we report in Table 7 the transport cost and FID values for samples from the marginals, for the 4 IMF iterations (note that we train a classifier and use the obtained features for the FID calculation, so these values should not be compared with those in other works). We initialise the sampling from 1000 unseen test samples of the digit 6, and report the FID values of the obtained 2s and 4s (averaged). We see that, as expected, the transport cost decreases as the IMF iterations proceed, indicating that the barycentre approximation is improving. We also observe that the FID scores increase slightly (though there is little visible difference) — this is a consequence of the drift that can accrue in the marginals, and is consistent with the expected behaviour of DSBM from which this effect is inherited (this can be reduced by training for longer or using the techniques discussed in Appendix C.1).
    Table 7: Progression of the total transport cost and fit to the marginals (as measured by FID), during the IMF iterations. Note that the FID values are obtained using a trained classifier, so should not be compared to values in other works.
    IMF 1 IMF 2 IMF 3 IMF 4
    Transport cost 431 402 389 378
    Ave. FID (2s and 4s) 61 82 89 93
@@ D.3 Subset posterior aggregation
    The previous experiments have shown TreeDSBM to improve over its IPF counterpart TreeDSB. In the following experiments, we provide a more detailed comparison with current strongly-performing methods for continuous Wasserstein-2 barycentre estimation, by reporting results for standard experiments in the literature. We include comparisons against the methods WIN [47], W2CB [49], and the recent method NOTWB [45]. These methods have demonstrated strong empirical performance in their respective works, and are chosen here to be representative of different approaches in the literature—WIN is an iterative method inspired by [4], W2CB is an Input Convex Neural Network-based approach [5, 62], and NOTWB is based on recent Neural OT methodology [50]. We use the implementations in their publicly available code, to which we provide the links in Appendix F. We note that, as ever in barycentre studies, it is somewhat challenging to assess the performance of solvers due to the lack of the ground-truth solution (other than in certain specific examples). Here, we report results on standard experiments used in the literature.
    Experimental setup
    We work with the experimental setup and dataset used in [49] (the same dataset was also used previously in [54, 28]), which uses Poisson and negative-binomial regressions on a bike-rental dataset [29]. The aim is to predict the hourly number of bike rentals using features including day of the week, weather conditions, and more. The dataset is 8-dimensional and is split into 5 distinct subsets each of size 100,000. The ‘ground-truth’ barycentre consists of 100,000 samples from the full dataset posterior.
    Following the literature, we report the $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ metric between the ‘ground-truth’ and the obtained samples in Table 3. For methods that generate from each marginal, we report the average over generations from each marginal, and for WIN we report results for the barycentre generator. The $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ metric is defined as
    where the Bures-Wasserstein metric is defined as $)\mathrm{BW}_{2}^{2}(\nu,\tilde{\nu})=W_{2}^{2}(\mathcal{N}(m_{\nu},\Sigma_{\nu}),\mathcal{N}(m_{\tilde{\nu}},\Sigma_{\tilde{\nu}}$)) for the respective means and covariances of the distributions.
    Hyperparameters
@@
    For W2CB, we use a learning rate of 1e-4 in the negative binomial setting. For WIN, in the negative binomial case we rescale the source $z$-sampler by a factor of 10.0 to match the scale of the data better; without this, it did not appear to converge. We run W2CB and WIN for 10000 training iterations, and NOTWB for 2500 iterations. Other than those mentioned, we use the default parameters provided in the respective codebases.
@@
    The W2CB algorithm appeared to give good results after approximately 1000 training steps in both cases, which took around 10 minutes in our experiments. WIN converged after around 2500 iterations, which took around 45 minutes. NOTWB converged quickly after only around 200 iterations, which took around 2 minutes.
    In both experiments, each TreeIMF iteration for our TreeDSBM implementation took approximately 20 seconds when training the edges jointly. TreeDSBM converged well using 4 IMF iterations, and training took around 2 minutes 30 seconds (including time for simulating training samples for the next iteration). This is comparable with NOTWB, the fastest of the alternative methods.
    Convergence speed
    We provide the values of the $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ metric as the IMF iterations progress in Table 8, and again we observe very fast convergence.
[…]
    IMF 1 IMF 2 IMF 3 IMF 4
    $\downarrow$ Poisson 31.1 $\pm$ 0.03 0.0085 $\pm$ 0.0003 0.0075 $\pm$ 0.0006 0.0076 $\pm$ 0.0005
    $\downarrow$ Negative Binomial 31.0 $\pm$ 0.01 0.0123 $\pm$ 0.0003 0.0118 $\pm$ 0.0007 0.0121 $\pm$ 0.0004
    D.4 Higher-dimensional Gaussian experiments
    Experimental setup
    We follow the experimental setup previously used in [49, 47, 45], in which 3 Gaussian distributions and its ground-truth ($\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3}$)-barycentre are randomly generated, for each dimension in $\{64,96,128\}$. We report results for $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ (which measures the quality of the overall generated barycentre) as described above, and additionally report the $L^{2}\mathrm{-UVP}$ metric, which measures the quality of the individual maps to the barycentre and is defined for each marginal as